# Prefix-free parsing for merging big BWTs

### v1

@@ 4 Experiments
We took from www.foodsafety.gov/food-poisoning/bacteria-and-viruses a list of 10 bacteria commonly causing food poisoning (botulism is caused by C. botulinum; norovirus and hepatitis A are viruses), searched for each one at www.ncbi.nlm.nih.gov/datasets and concatenated the FASTA listed on the first page of results. The resulting 10 datasets were between 73 MB and 237 MB, with a total size of 1.71 GB; and between 74 MB and 234 MB, with a total size of 1.69 GB when we removed newlines and replaced other non-$\{\mathtt{A},\mathtt{C},\mathtt{G},\mathtt{T}\}$ characters by copies of X.
We built the BWT or eBWT for the collection of datasets using Manzini’s bigbwt [^1] https://gitlab.com/manzai/Big-BWT, which uses standard PFP; Li’s ropebwt3222https://github.com/lh3/ropebwt3; and an implementation of our algorithm, PFP-merge [^3] We will post our code when this paper is accepted.. According to time -v, on a commodity laptop with an Intel® Core™ i5-1335U $\times$ 12 processor and 16 GB of RAM running Ubuntu 24.04.2 LTS
  • bigbwt on the concatenated filtered files with default parameters took 354.64 seconds of user time, 5.12 seconds of system time, 360.20 seconds of wall-clock time, and 4 529 976 KB peak memory;
  • ropebwt3 on the separate FASTA files with default parameters (but with the -R flag to prevent it indexing its input’s reversed complement as well) took 3034.65 seconds of user time, 21.45 seconds of system time, 345.34 seconds of wall-clock time, and 3 720 816 KB peak memory;
  • ropebwt3 on the concatenation of the FASTA files with default parameters (but with -R) ran out of memory;
  • PFP-merge on the separate filtered files with default parameters ($w=20$, $p=100$) took 549.35 seconds of user time, 10.21 seconds of system time, 560.15 seconds of wall-clock time, and 1 004 972 KB peak memory.
In short, although PFP-merge took 1.56 and 1.62 times as much wall-clock time as bigbwt and ropebwt3, respecctively, it took only 0.22 and 0.27 times as much peak memory. Moreover, the actual merging took only 205.71 seconds and 287 104 KB of peak memory — slightly more than 3/8ths the total size of the small datasets’ dictionaries (630 307 KB), as we expected — suggesting PFP-merge will do relatively better on more small datasets.
@@ 4.0.1 Acknowledgements
Many thanks to Lavinia Egidi, Felipe Louza and Giovanna Rosone for helpful discussions, and belated thanks to Paweł Gawrychowski, Dmitry Kosolobov and Tatiana Starikovskaya for suggestions that sped up PFP in the past. TG funded by NSERC grant RGPIN-07185-2020. BL and VS funded by NIH grant R01HG011392 to BL. ZsL partially funded by MUR PRIN project no. 2022YRB97K and INdAM-GNCS project no. E53C24001950001.

### v2

@@ 4 Experiments
We downloaded genomes of 30 species of bacteria from the AllTheBacteria project [^1] https://allthebacteria.org/, with sizes ranging from 309.93 MB to 314.53 MB and totalling 9.39 GB. We built their BWTs or eBWTs with our pfp-merge [^2] https://gitlab.com/manzai/pfp-merge, Manzini’s bigbwt [^3] https://gitlab.com/manzai/Big-BWT (based on plain PFP), Díaz’s grlbwt [^4] https://github.com/ddiazdom/grlBWT, Li’s ropebwt3555https://github.com/lh3/ropebwt3 and Olbrich’s lg [^6] https://gitlab.com/qwerzuiop/lyndongrammar (for “Lyndon grammar”). We omitted Masillo’s CMT-BWT because it requires a single reference. We used default parameter settings for all the programs (except we used ropebwt3’s -R flag to prevent it indexing also the datasets’ reverse complements) and, due to time constraints, 16 threads for each program on similar nodes of a cluster at the University of Helsinki. For pfp-merge, bigbwt, and lg, we removed newlines and converted non-$\{\mathtt{A},\mathtt{C},\mathtt{G},\mathtt{T}\}$ characters to Xs. We ran ropebwt3 on the original FASTA files and grlbwt on the concatenation of the datasets (one string per line). Table 1 shows the wall-clock time and peak memory used by each program. Although pfp-merge was almost the slowest — which may improve with additional parallelization — it used the least memory by a noticeable margin. The actual merging used only 0.77 GB of memory, suggesting we may perform relatively even better on more datasets.
Table 1: The time and peak memory used by each program on 30 bacterial pangenomes totalling 9.39 GB.
time memory
(mm:ss) (GB)
pfp-merge 23:50 0.91
bigbwt 16:26 13.84
glrbwt 9:01 1.71
ropebwt3 27:10 5.12
lg 7:41 7.00
@@ 4.0.1 Acknowledgements
Many thanks to Lavinia Egidi, Felipe Louza and Giovanna Rosone for helpful discussions, and belated thanks to Paweł Gawrychowski, Dmitry Kosolobov and Tatiana Starikovskaya for suggestions that sped up PFP in the past. TG funded by NSERC grant RGPIN-07185-2020. BL and VS funded by NIH grant R01HG011392 to BL. ZsL partially funded by MUR PRIN project no. 2022YRB97K and INdAM-GNCS project no. E53C24001950001. VG and GM funded by the Next Generation EU PNRR MUR M4 C2 Inv 1.5 project ECS00000017 Tuscany Health Ecosystem Spoke 6 CUP I53C22000780001.